# TRIM: Token-wise Attention-Derived Saliency for Data-Efficient Instruction Tuning

### v1

@@ 6 Conclusion
We introduced TRIM, a token-centric framework for efficient coreset selection in instruction tuning. By shifting from coarse, sample-level signals to fine-grained, attention-derived token representations, TRIM addresses two persistent challenges in data selection: computational cost and length bias. Through saliency-weighted fingerprints constructed from a handful of target samples, TRIM identifies training data that matches the structural patterns defining a task, rather than relying on expensive gradient computation or surface-level similarity. Our experiments demonstrate that TRIM consistently outperforms state-of-the-art methods across several downstream tasks, achieving up to 9% improvements with only 5% coresets. Notably, TRIM’s token-level scoring enables it to surface structurally relevant data even in low-overlap settings, as evidenced by near-full-data performance on mathematical reasoning using a general-purpose corpus. The method’s forward-only design delivers these gains at a fraction of the computational cost of gradient-based alternatives, and coresets selected by small scorers transfer effectively to larger models across architectures. TRIM’s ability to capture task-defining structure through attention signals offers a scalable path forward, enabling practitioners to curate smaller, higher-quality datasets.
Limitations
TRIM’s effectiveness depends on the quality and representativeness of the target validation set. With only 5-10 examples guiding selection, noise or bias in these samples can propagate into the fingerprints, potentially degrading coreset quality. Future work could explore ensemble approaches that aggregate fingerprints across multiple validation sets to improve robustness. Additionally, TRIM requires a target set, making it not directly compatible for pretraining data curation. However, the core principle of matching data-model interactions can be potentially extended to pretraining through task-agnostic quality signals, for instance, using broad capability benchmarks as proxy validation sets.
@@ Appendix A Extended Related Work
This section provides extended details on prior work in data selection for instruction tuning, expanding upon the summary in section˜2.
@@
Influence-based methods aim to quantify the causal effect of a candidate training example on model performance with respect to a target task. A foundational approach, Influence Functions (Koh and Liang, 2017), approximates the impact of upweighting a training point via second-order Hessian computations. As these are infeasible for LLMs, scalable surrogates have been proposed. DataInf (Kwon et al., 2024) and In2Core (San Joaquin et al., 2024) exploit the low-rank structure of LoRA finetuning to derive closed-form influence approximations, enabling per-sample scoring. DealREC (Lin et al., 2024a) instead estimates Hessian-vector products on surrogate models and adjusts with gradient norms as proxies for learning difficulty.
@@
This family includes efficient, task-agnostic methods. BM25 (Robertson et al., 2009) and DSIR (Xie et al., 2023) select based on lexical overlap or $n$-gram distributions, but lack target awareness. Recent works employ lightweight scorers or LLMs-as-judges to filter or rank examples. Examples include filtering low-quality responses (Chen et al., 2024), ranking samples for diversity (Ge et al., 2024), or targeting difficult instructions (Li et al., 2024; Liu et al., 2024a). These methods are efficient but depend on proxy notions of data quality.
@@
A principled line of work casts data selection as optimization. TSDS (Liu et al., 2024b) minimizes distribution alignment loss between coresets and target distributions. GREATS (Wang et al., 2024) greedily selects data by gradient alignment with validation batches. PDS (Gu et al., 2025) frames the task as optimal control, linking selection scores to downstream impact. These methods are elegant but computationally heavy.
@@
Recognizing the limitations of sample-level metrics, the most recent work has begun to shift towards token-level signals. These methods leverage the internal mechanisms of Transformers to find important data. For instance, LADM uses attention scores to measure long-range dependencies for long-context data selection (Chen et al., 2025). T-SHIRT proposes a hierarchical filtering scheme for general-purpose data pruning, rather than for the task-specific, targeted selection that we address (Fu et al., 2025). Its complex pipeline first uses K-Means clustering for a coarse, document-level selection, then refines this by scoring individual tokens via iterative input perturbations, making it computationally intensive.
@@
Furthermore, our multi-layer approach to calculating salience is inspired by interpretability research. Work on "Quantifying Attention Flow" (Abnar and Zuidema, 2020) established the value of aggregating attention across all layers to build a complete information flow graph, thereby determining a token’s overall importance. While their graph-based max-flow algorithm is designed for deep interpretability, TRIM adapts the underlying principle into a fast heuristic (focus and centrality) suitable for a large-scale, practical data selection task.
@@ Appendix B Pseudocode for TRIM
1:Warmed model $\mathbf{M}_{\mathrm{warmup}}$, target val $\mathcal{T}_{\mathrm{val}}$, Number of layers $L$
2:Salience weights $(w_{Q},w_{K})=(\tfrac{1}{2},\tfrac{1}{2})$
3:Fingerprint dictionary $\mathcal{F}=\{f_{t}\}$
[…]
5:for each $v\in\mathcal{T}_{\mathrm{val}}$ do
6:  Run $\mathbf{M}_{\mathrm{warmup}}$ once (attn+states, last $L$)
7:  for each position $i$ in $v$ do
8:   Compute $Q_{i}$ and $\tilde{K}_{i}$ (eqs. in text)
9:   $\alpha_{i}\leftarrow w_{Q}Q_{i}+w_{K}\tilde{K}_{i}$
10:for each token type $t$ seen in $\mathcal{T}_{\mathrm{val}}$ do
11:  $O_{t}\leftarrow\{(v,i):\mathrm{type}(v_{i})=t\}$
12:  $\hat{h}_{v,i}\leftarrow h_{v,i}/\|h_{v,i}\|_{2}$
13:  $f_{t}\leftarrow\mathrm{normalize}\!\left(\sum_{(v,i)\in O_{t}}\alpha_{v,i}\hat{h}_{v,i}\right)$
14:  $\mathcal{F}[t]\leftarrow f_{t}$
@@
[…]
2:Pool weights $(w_{\mu},w_{m})=(\tfrac{1}{2},\tfrac{1}{2})$, Non fingerprinted penalty $\lambda\in(0,1)$
3:Scores $\{S(c)\}_{c\in\mathcal{S}}$
4:for each $c\in\mathcal{S}$ do
5:  Run $\mathbf{M}_{\mathrm{warmup}}$ once to get $\hat{h}_{c,j}$ for all $j$
6:  $M(c)\leftarrow\emptyset$
7:  for each position $j$ do
8:   $t_{j}\leftarrow\mathrm{type}(c_{j})$
9:   if $t_{j}\in\mathcal{F}$ then
10:     $s_{j}\leftarrow\cos(\hat{h}_{c,j},\,\mathcal{F}[t_{j}])$
11:   else
12:     $\bar{t}_{j}\leftarrow\arg\max_{t\in\mathcal{F}}\cos(e_{t_{j}},e_{t})$
13:     $s_{j}\leftarrow\lambda\cdot\cos(\hat{h}_{c,j},\,\mathcal{F}[\bar{t}_{j}])$
14:   $M(c)\leftarrow M(c)\cup\{j\}$
15:  if $|M(c)|=0$ then
16:   $S(c)\leftarrow$-$\infty$
17:  else
18:   $\mu\leftarrow\frac{1}{|M(c)|}\sum_{j\in M(c)}s_{j}$
19:   $m\leftarrow\max_{j\in M(c)}s_{j}$
20:   $S(c)\leftarrow w_{\mu}\mu+w_{m}m$
21:return $\{S(c)\}$
Algorithm 3 TRIM Framework
(Warmup $\rightarrow$ Stage I $\rightarrow$ Stage II $\rightarrow$ Selection)
1:Base LLM $\mathbf{M}_{0}$, pool $\mathcal{S}$, target val $\mathcal{T}_{\mathrm{val}}$, coreset size $K$
2:Coreset $\mathcal{C}\subset\mathcal{S}$ of size $K$
3:Warmup: sample $\mathcal{S}_{\mathrm{warmup}}\subset\mathcal{S}$ ($\sim$5%); fine-tune $\mathbf{M}_{0}$
[…]
8:return $\mathcal{C}$
We provide detailed pseudocode for the complete TRIM pipeline to complement the description in Section˜3.
The main driver, presented in Algorithm˜3, orchestrates the overall workflow: model warmup, fingerprint generation, candidate scoring, and final coreset selection. This algorithm relies on two core subroutines that correspond to the two stages of our method.
Stage I, implemented in BuildFingerprints (Algorithm˜1), details the logic from Section˜3.1. It processes the target validation set $\mathcal{T}_{\text{val}}$ using the warmed-up model $\mathbf{M}_{\text{warmup}}$. For each token, it computes the attention-derived salience $\alpha_{i}$ by combining the director ($Q_{i}$) and hub ($\tilde{K}_{i}$) scores. These salience values are then used to create the final, salience-weighted token fingerprints $\mathcal{F}=\{f_{t}\}$.
Stage II, detailed in ScoreCandidates (Algorithm˜2), executes the scoring process from Section˜3.2. For each candidate, it computes token-wise similarity to the fingerprints, applying a penalized backoff for out-of-vocabulary types. These token scores are then aggregated into a robust example score using a mean-max pooling strategy with a coverage bonus.
@@
B.1 Nuances in the Algorithm
@@
If selective token scoring is done, then the token at position $j$ is included in $M(c)$ if and only if it falls within the chosen scoring scope (prompt, response, or all) and is not a special token (e.g., BOS, EOS, PAD). The token’s score $s_{j}$ is calculated as defined above, whether through a direct fingerprint match or OOV backoff.
$M(c)=\{\,j:\ \text{token }j\text{ is scored}\,\}.$ (14)
If, for a given candidate, this set is empty $\left(|M(c)|=0\right)$, we assign it a score of $S(c)=$-$\infty$. And the robust pooling operation will then include an additional penalty term called “coverage" ($\kappa(c)$).
$\mu(c)\;=\;$ $\frac{1}{|M(c)|}\sum_{j\in M(c)}s_{j},$
$m(c)\;=\;$ $\max_{j\in M(c)}s_{j},$
$\kappa(c)\;=\;$ $\frac{|M(c)|}{|c|}.$
$S(c)\;=\;w_{\mu}\,\mu(c)\;+\;w_{m}\,m(c)\;+\;\eta\,\kappa(c).$ (15)
@@
Appendix C Baseline Data Selection Strategy
@@
Low-rank gradient Similarity Search (LESS) is an optimizer-aware method for targeted data selection. It adapts the classic gradient-similarity influence formulation to work with the Adam optimizer and variable-length instruction data. To remain efficient, LESS uses LoRA and random projections to compute a low-dimensional "gradient datastore" from a short warmup training phase. Candidates are scored by the cosine similarity of their gradient features to those of the target validation set, aggregated across several checkpoints.
@@
Appendix D Training Datasets Overview

### v2

@@ 6 Conclusion
We introduced TRIM, a token-centric framework for efficient coreset selection in instruction tuning. By shifting from coarse, sample-level signals to fine-grained, attention-derived token representations, TRIM addresses two persistent challenges in data selection: computational cost and length bias. Through saliency-weighted fingerprints constructed from a handful of target samples, TRIM identifies training data that matches the structural patterns defining a task, rather than relying on expensive gradient computation or surface-level similarity. Our experiments demonstrate that TRIM consistently outperforms state-of-the-art methods across several downstream tasks, achieving up to 9% improvements with only 5% coresets. Notably, TRIM’s token-level scoring enables it to surface structurally relevant data even in low-overlap settings, as evidenced by near-full-data performance on mathematical reasoning using a general-purpose corpus. We also show that TRIM extends to setting where there are no separate target examples available, enabling effective curation directly from a specialized training pool. The method’s forward-only design delivers these gains at a fraction of the computational cost of gradient-based alternatives, and coresets selected by small scorers transfer effectively to larger models across architectures. TRIM’s ability to capture task-defining structure through attention signals offers a scalable path forward, enabling practitioners to curate smaller, higher-quality datasets.
Impact Statement
TRIM improves the computational efficiency of instruction-tuning by enabling targeted coreset selection using forward-only signals and token-level structure matching. By reducing the amount of training data (and associated optimization steps) needed to achieve strong downstream performance, TRIM can lower the energy and cost of adapting language models, potentially broadening access to model customization for academic groups and smaller organizations. At the same time, TRIM is a data curation method: it changes which examples are used for training. If the candidate pool contains harmful, biased, or sensitive content, selection may amplify those properties by prioritizing examples that align with a target reference set or with in-domain structural patterns. In particular, targeted selection can overemphasize narrow behaviors present in the reference samples, which may reduce diversity, increase spurious correlations, or degrade performance on underrepresented groups and topics. As a result, practitioners should treat coreset selection as part of the safety pipeline, auditing candidate pools, filtering toxic or private data, and evaluating for fairness and robustness after selection.
TRIM is not designed to detect or remove unsafe content, nor does it provide guarantees about privacy or bias. The method also does not prevent misuse of instruction tuning to produce models optimized for harmful objectives; it only makes selection more efficient. We recommend using TRIM alongside established dataset governance practices (licensing checks, PII filtering, toxicity screening) and downstream evaluations that include safety-focused benchmarks and red-teaming where appropriate.
Overall, we expect TRIM’s primary impact to be enabling more compute-efficient and targeted dataset curation for model adaptation, while emphasizing that responsible deployment requires careful dataset auditing and post-selection evaluation to mitigate risks of bias amplification and harmful content retention.
@@ Appendix A Extended Related Work
This section provides extended details on prior work in data selection for instruction tuning, expanding upon the summary in Section 2.
@@
Influence-based methods aim to quantify the causal effect of a candidate training example on model performance with respect to a target task. A foundational approach, Influence Functions (Koh and Liang, 2017), approximates the impact of upweighting a training point via second-order Hessian computations. As these are infeasible for LLMs, scalable surrogates have been proposed. DataInf (Kwon et al., 2024) and In2Core (San Joaquin et al., 2024) exploit the low-rank structure of LoRA finetuning to derive closed-form influence approximations, enabling per-sample scoring. DealREC (Lin et al., 2024b) instead estimates Hessian-vector products on surrogate models and adjusts with gradient norms as proxies for learning difficulty.
@@
This family includes efficient, task-agnostic methods. BM25 (Robertson et al., 2009) and DSIR (Xie et al., 2023) select based on lexical overlap or $n$-gram distributions, but lack target awareness. Recent works employ lightweight scorers or LLMs-as-judges to filter or rank examples. Examples include filtering low-quality responses (Chen et al., 2024a), ranking samples for diversity (Ge et al., 2024), or targeting difficult instructions (Li et al., 2024; Liu et al., 2024a). These methods are efficient but depend on proxy notions of data quality.
@@
A principled line of work casts data selection as optimization. TSDS (Liu et al., 2024b) minimizes distribution alignment loss between coresets and target distributions. GREATS (Wang et al., 2024) greedily selects data by gradient alignment with validation batches. PDS (Gu et al., 2025) frames the task as optimal control, linking selection scores to downstream impact. More recently, QCS (Chen et al., 2025b) formulates coreset construction as a bi-level optimization problem that couples sequence-level selection with token-level mining. In principle, differentiating through the lower-level token mining step can introduce second-order terms (e.g., Hessian-related hypergradients), and QCS proposes a relaxed probabilistic variant to avoid explicit Hessian computation. Overall, while these approaches are conceptually elegant, their optimization loops (and associated higher-order machinery or relaxations) can remain computationally heavy at scale.
@@
Recognizing the limitations of sample-level metrics, the most recent work has begun to shift towards token-level signals. These methods leverage the internal mechanisms of Transformers to find important data. For instance, LADM uses attention scores to measure long-range dependencies for long-context data selection (Chen et al., 2025a). T-SHIRT proposes a hierarchical filtering scheme for general-purpose data pruning, rather than for the task-specific, targeted selection that we address (Fu et al., 2025). Its complex pipeline first uses K-Means clustering for a coarse, document-level selection, then refines this by scoring individual tokens via iterative input perturbations, making it computationally intensive.
@@
Furthermore, our multi-layer approach to calculating salience is inspired by interpretability research. Work on ”Quantifying Attention Flow” (Abnar and Zuidema, 2020) established the value of aggregating attention across all layers to build a complete information flow graph, thereby determining a token’s overall importance. While their graph-based max-flow algorithm is designed for deep interpretability, TRIM adapts the underlying principle into a fast heuristic (focus and centrality) suitable for a large-scale, practical data selection task.
@@ Appendix B Pseudocode for TRIM
0:  Warmed model $\mathbf{M}_{\mathrm{warmup}}$, target val $\mathcal{T}_{\mathrm{val}}$, number of layers $L$
0:  Salience weights $(w_{Q},w_{K})=(\tfrac{1}{2},\tfrac{1}{2})$
0:  Fingerprint dictionary $\mathcal{F}=\{f_{t}\}$
1:  $\mathcal{F}\leftarrow\emptyset$
2:  for each $v\in\mathcal{T}_{\mathrm{val}}$ do
3:  Run $\mathbf{M}_{\mathrm{warmup}}$ once (attn+states, last $L$)
4:  for each position $i$ in $v$ do
5:   Compute $Q_{i}$ and $\tilde{K}_{i}$ (eqs. in text)
6:   $\alpha_{i}\leftarrow w_{Q}Q_{i}+w_{K}\tilde{K}_{i}$
7:  end for
8:  end for
9:  for each token type $t$ seen in $\mathcal{T}_{\mathrm{val}}$ do
10:  $O_{t}\leftarrow\{(v,i):\mathrm{type}(v_{i})=t\}$
11:  $\hat{h}_{v,i}\leftarrow h_{v,i}/\|h_{v,i}\|_{2}$
12:  $f_{t}\leftarrow\mathrm{normalize}\!\left(\sum_{(v,i)\in O_{t}}\alpha_{v,i}\hat{h}_{v,i}\right)$
13:  $\mathcal{F}[t]\leftarrow f_{t}$
14:  end for
@@
0:  $\mathbf{M}_{\mathrm{warmup}}$, pool $\mathcal{S}$, fingerprints $\mathcal{F}$
0:  Pool weights $(w_{\mu},w_{m})=(\tfrac{1}{2},\tfrac{1}{2})$, nearest-fingerprint fallback penalty $\lambda\in(0,1)$
0:  Scores $\{S(c)\}_{c\in\mathcal{S}}$
1:  for each $c\in\mathcal{S}$ do
2:  Run $\mathbf{M}_{\mathrm{warmup}}$ once to get $\hat{h}_{c,j}$ for all $j$
3:  $\mathcal{J}(c)\leftarrow\emptyset$
4:  for each position $j$ do
5:   $t_{j}\leftarrow\mathrm{type}(c_{j})$
6:   if $t_{j}\in\mathcal{F}$ then
7:    $s_{j}\leftarrow\cos(\hat{h}_{c,j},\,\mathcal{F}[t_{j}])$
8:   else
9:    $\bar{t}_{j}\leftarrow\arg\max_{t\in\mathcal{F}}\cos(e_{t_{j}},e_{t})$
10:    $s_{j}\leftarrow\lambda\cdot\cos(\hat{h}_{c,j},\,\mathcal{F}[\bar{t}_{j}])$
11:   end if
12:   $\mathcal{J}(c)\leftarrow\mathcal{J}(c)\cup\{j\}$
13:  end for
14:  if $|\mathcal{J}(c)|=0$ then
15:   $S(c)\leftarrow$-$\infty$
16:  else
17:   $\mu\leftarrow\frac{1}{|\mathcal{J}(c)|}\sum_{j\in\mathcal{J}(c)}s_{j}$
18:   $m\leftarrow\max_{j\in\mathcal{J}(c)}s_{j}$
19:   $S(c)\leftarrow w_{\mu}\mu+w_{m}m$
20:  end if
21:  end for
22:  return $\{S(c)\}$
Algorithm 3 TRIM Framework (Warmup $\rightarrow$ Stage I $\rightarrow$ Stage II $\rightarrow$ Selection)
0:  Base LLM $\mathbf{M}_{0}$, pool $\mathcal{S}$, target val $\mathcal{T}_{\mathrm{val}}$, coreset size $K$
0:  Coreset $\mathcal{C}\subset\mathcal{S}$ of size $K$
1:  Warmup: sample $\mathcal{S}_{\mathrm{warmup}}\subset\mathcal{S}$ ($\sim$5%); fine-tune $\mathbf{M}_{0}$
2:  $\mathbf{M}_{\mathrm{warmup}}\leftarrow$ adapted model
3:  $\mathcal{F}\leftarrow\textsc{BuildFingerprints}(\mathbf{M}_{\mathrm{warmup}},\mathcal{T}_{\mathrm{val}},L)$
4:  $\{S(c)\}\leftarrow\textsc{ScoreCandidates}(\mathbf{M}_{\mathrm{warmup}},\mathcal{S},\mathcal{F})$
5:  $\mathcal{C}\leftarrow$ top-$K$ candidates by $S(c)$
6:  return $\mathcal{C}$
We provide detailed pseudocode for the complete TRIM pipeline to complement the description in Section 3. The main driver, presented in Algorithm 3, orchestrates the overall workflow: model warmup, fingerprint generation, candidate scoring, and final coreset selection. This algorithm relies on two core subroutines that correspond to the two stages of our method.
Stage I, implemented in BuildFingerprints (Algorithm 1), details the logic from Section 3.1. It processes the target validation set $\mathcal{T}_{\text{val}}$ using the warmed-up model $\mathbf{M}_{\text{warmup}}$. For each token, it computes the attention-derived salience $\alpha_{i}$ by combining the director ($Q_{i}$) and hub ($\tilde{K}_{i}$) scores. These salience values are then used to create the final, salience-weighted token fingerprints $\mathcal{F}=\{f_{t}\}$.
Stage II, detailed in ScoreCandidates (Algorithm 2), executes the scoring process from Section 3.2. For each candidate, it computes token-wise similarity to the fingerprints, applying a penalized backoff for out-of-vocabulary types. These token scores are then aggregated into a robust example score using a mean-max pooling strategy with a coverage bonus.
@@
Appendix C Details in the Algorithm
Handling Non-fingerprinted Classes
The Nearest fingerprint fallback (NFF) mapping is designed to handle token classes that do not appear in the target set (and hence lack fingerprints) without introducing per-example nearest-neighbor overhead. Let $\mathcal{V}$ be the vocabulary and $\mathcal{F}\subseteq\mathcal{V}$ the fingerprinted token classes. We unit-normalize all input embeddings $\{e_{t}\}_{t\in\mathcal{V}}$ from the warmup model and precompute a lookup table
$\pi:\mathcal{V}\rightarrow\mathcal{F}\quad\text{where}\quad\pi(t)=\arg\max_{u\in\mathcal{F}}\cos(e_{t},e_{u})$
This precomputation is done once per tokenizer/model (offline) and cached; at scoring time, each non-fingerprinted token $t_{j}$ incurs only an $O(1)$ lookup to obtain $\bar{t}_{j}=\pi(t_{j})$ and then uses the penalized similarity $s_{j}=\lambda\cdot\cos(\hat{h}_{c,j},f_{\bar{t}_{j}})$.
In practice, this caching makes NFF negligible in the overall selection runtime, and it ensures that unseen token classes contribute conservatively via the penalty $\lambda$ rather than being discarded entirely.
@@ Appendix C Details in the Algorithm
If selective token scoring is done, then the token at position $j$ is included in $\mathcal{J}(c)$ if and only if it falls within the chosen scoring scope (prompt, response, or all) and is not a special token (e.g., BOS, EOS, PAD). The token’s score $s_{j}$ is calculated as defined above, whether through a direct fingerprint match or NFF backoff.
$\mathcal{J}(c)=\{\,j:\ \text{token }j\text{ is scored}\,\}.$ (14)
If, for a given candidate, this set is empty $\left(|\mathcal{J}(c)|=0\right)$, we assign it a score of $S(c)=$-$\infty$. And the robust pooling operation will then include an additional penalty term called “coverage” ($\kappa(c)$).
$\mu(c)\;=\;$ $\frac{1}{|\mathcal{J}(c)|}\sum_{j\in\mathcal{J}(c)}s_{j},$
$\tau(c)\;=\;$ $\max_{j\in\mathcal{J}(c)}s_{j},$
$\kappa(c)\;=\;$ $\frac{|\mathcal{J}(c)|}{|c|}.$
$S(c)\;=\;w_{\mu}\,\mu(c)\;+\;w_{m}\,\tau(c)\;+\;\eta\,\kappa(c).$ (15)
@@
IDF-scaled fingerprinting.
TRIM can optionally incorporate inverse document frequency (IDF) to reduce the influence of ubiquitous tokens when constructing fingerprints. Concretely, for a token class $t$, we scale each occurrence weight by $\mathrm{IDF}(t)$ (computed over the candidate pool), i.e., we replace $\alpha_{v,i}$ with $\alpha_{v,i}\cdot\mathrm{IDF}(t)$ in Eq. (9). This can be helpful when the candidate pool contains substantial shared formatting or “template” text (which otherwise yields many high-saliency but low-specificity matches), while having negligible impact when the task signal is already distributed across many diverse tokens (e.g., MMLU; see Table 11).
Appendix D Baseline Data Selection Strategies
@@
Low-rank gradient Similarity Search (LESS) is an optimizer-aware method for targeted data selection. It adapts the classic gradient-similarity influence formulation to work with the Adam optimizer and variable-length instruction data. To remain efficient, LESS uses LoRA and random projections to compute a low-dimensional “gradient datastore” from a short warmup training phase. Candidates are scored by the cosine similarity of their gradient features to those of the target validation set, aggregated across several checkpoints.
@@
Appendix E Training Datasets Overview